OpenStudy (anonymous):

Given a = -3 and b = 4 and c = -5, evaluate |c - a - b|.

OpenStudy (anonymous):

Will you STOP spamming? Jesus.

OpenStudy (anonymous):

l-5-(-3)-4l = l-2-4l = l-6l =6

OpenStudy (shadowfiend):

odubo, rather than reporting malevolence19 for abuse, please stop spamming.

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

Thank you shadow.

OpenStudy (anonymous):

it would be 6 still because the two lines around the number are for absolute value, which is how far it is from zero. so 6 is correct.